### v1

@@ Abstract
Vision–language models (VLMs) have shown strong promise for medical image analysis, but most remain opaque, offering predictions without the transparent, stepwise reasoning clinicians rely on. We present a framework that brings chain-of-thought (CoT) reasoning to chest X-ray interpretation, and release the model NV-Reason-CXR-3B.  [^1] https://github.com/NVIDIA-Medtech/NV-Reason-CXR [^2] https://huggingface.co/nvidia/NV-Reason-CXR-3B
Inspired by reasoning-first training paradigms, our approach is designed to learn how experts reason—not just what they conclude—by aligning intermediate steps with observable image evidence and radiology workflow. Beyond accuracy, the explicit reasoning traces support clinical auditability: they reveal why a conclusion was reached, which alternatives were considered, and where uncertainty remains—enabling quality assurance, error analysis, and safer human–AI collaboration. The model name is
Our model couples high-fidelity visual encoding with a two-stage training recipe: a reasoning-style supervised finetunning (SFT) followed by reinforcement learning (RL) that uses verifiable rewards over a list of X-ray abnormalities. The model outputs reasoning that mirrors radiologists’ systematic though process, uncertainty and differential diagnosis. In out-of-distribution evaluation, the approach achieves competitive multi-label classification while improving interpretability. In a reader study with expert radiologists, full reasoning traces increased confidence, supported error auditing, and reduced time to finalize reports. We release code and model to facilitate community progress toward trustworthy, explainable AI in chest radiography and broader medical imaging tasks where reasoning quality is as critical as prediction quality.
@@ 2.1 Human annotations
We used the OpenAI’s whisper-large-v3 model to transcribe audio to text [7]. The system transcribes in the background; radiologists then review the transcript and edit if necessary. We observed frequent transcription errors, particularly for medical terminology. To mitigate this, we added a second pass using an LLM for domain-aware correction. We instructed the LLM that the text concerns chest X-ray radiology to bias corrections toward appropriate terms. This substantially reduced manual edits by radiologists, who previously had to fix spelling and terminology errors. For this correction stage we used the DeepSeek-V3-0324 LLM model [8].
@@ 2.4 Multi-Turn Follow-Up Chat Data
Given the per-case follow-up QA pool, we constructed a multi-turn chat dataset where each dialogue sampled a random number (0–5) of follow-up turns. We also injected queries from the refusal set to simulate incorrect or off-topic follow-ups. The first user turn was always a comprehensive request (e.g., “Analyze the chest X-ray” or “Find abnormalities and support devices”) for two reasons: (a) to obtain a fully formatted reasoning answer that can later be used for RL training, and (b) to make subsequent follow-up questions easier to answer by establishing a systematic, context-rich baseline.
@@ 2.5 Structured Report Query
Structured report. INDICATION: Not provided. COMPARISON: None. TECHNIQUE: PA view of the chest. FINDINGS: Medical devices: Chest tube at right lung base. Central airways: Trachea and main bronchi patent. Lungs and pleural space: Small right pneumothorax; small right pleural effusion; parenchymal opacity in right lung; left lung clear. Mediastinum and hila: Mediastinum normal; hila normal; no adenopathy. Heart: Enlarged heart. Upper abdomen and diaphragm: No free subdiaphragmatic air; diaphragm normal. Bones and soft tissues: Normal. IMPRESSION: Small right-sided pneumothorax. Small right pleural effusion with parenchymal opacity. Enlarged heart
@@ 3.5 Verifiable Rewards for Chest X-Rays
The main reward function we use is abnormality set correctness, which is a intersection over union metric between predicted and ground-truth sets. Let $\hat{z}_{\text{cls}}=\Phi_{\text{cls}}(y)$ be the set extracted from $y$.
@@
Weights can upweight safety-critical findings (e.g., Lung Lesion), but in practice we used equal weights $w_{a}==1$. This reward function is by definition bound to 0..1 interval.
@@ 3.6 Training Details
Backbone. We adopt Qwen2.5-VL-3B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct, a vision–language model that combines a ViT-based vision encoder with a decoder-only LLM. Qwen2.5-VL features several architectural and training improvements, e.g. an efficient vision stack (windowed attention in ViT) and alignment to the language backbone with RMSNorm and SwiGLU—which together yield strong zero-/few-shot multimodal performance while remaining parameter-efficient for domain adaptation [15].
@@ 3.7 SFT
The training was performed on 4 nodes of 8xH100 NVIDIA GPU machines. All experiments use $bfloat16$ training with gradient checkpointing, gradient accumulation steps of 8. For GRPO each prompt has number of generations equal to 16, with temperature of 1. Each gpu was set to handle 1 prompt. AdamW optimizer was used with cosine schedule starting from learning rate of $1e^{-6}$. Grad norm clipping was set to 0.2. Optimization was run for 10 epochs.
@@ 3.9 Implementation details
Many design choices were inspired by Open-R1444https://github.com/huggingface/open-r1 - a great open source project by Hugginface aiming to reproduce DeepSeek-R1 [2] performance for text only inputs in PyTorch. Both SFT and GRPO training loops take advantage of several Hugginface libraries, including TRL, Transformers and Accelerate. We make our training code public [^5] https://github.com/NVIDIA-Medtech/NV-Reason-CXR.
@@ 4.1 Subjective Evaluation with Expert Radiologists
    After a short distractor, the case is re-presented with the abnormalities-only list (AI generated); the reader may revise findings/impression and time.
@@
    After a second distractor, the case is shown with the full reasoning text and its structured report form. The reader can copy-paste the structured report, double check it and edit if necessary. The total time is recorded again.
@@ 4.2 Numeric Evaluation
To evaluate multi-label prediction, model outputs are taken from the structured <answer> block which are already mapped to 14 classes. MedGemini work considered both multi-select prompt (asking for all labels at once) and per-label Yes/No prompts, and found per-label prompts provided better macro-F1 metric scores. In our work we use a single multi-select prompt to be consistent with the overall model workflow. The prompt was ”Find abnormalities and support devices”, and the reasoning text in <think> tags was ignored.
@@
Training used data only from the MIMIC-CXR training set. The SFT stage used a subset of 100K cases (images + reports), whereas RL stage used all 227K cases (images + report derived labels). OOD evaluation uses the official CheXpert test set.
@@
We report macro F1 score. Clinical classification tasks are often class-imbalanced: positive (abnormal) cases are relatively rare, yet clinically critical. Metrics that are dominated by the majority class (e.g., accuracy) can therefore be misleading. The F1 score directly balances precision and recall (sensitivity), making it well-suited to summarize performance when both missed findings and false alarms have tangible clinical costs.

### v2

@@ Abstract
Vision–language models (VLMs) have shown strong promise for medical image analysis, but most remain opaque, offering predictions without the transparent, stepwise reasoning clinicians rely on. We present a framework that brings chain-of-thought (CoT) reasoning to chest X-ray interpretation. Inspired by reasoning-first training paradigms, our approach is designed to learn how experts reason—not just what they conclude—by aligning intermediate steps with observable image evidence and radiology workflow. Beyond accuracy, the explicit reasoning traces support clinical auditability: they reveal why a conclusion was reached, which alternatives were considered, and where uncertainty remains—enabling quality assurance, error analysis, and safer human–AI collaboration.
Our model couples high-fidelity visual encoding with a two-stage training recipe: a reasoning-style supervised fine-tuning (SFT) followed by reinforcement learning (RL) that uses verifiable rewards over a list of X-ray abnormalities. The model outputs reasoning that mirrors radiologists’ systematic though process, uncertainty and differential diagnosis. In out-of-distribution evaluation, the approach achieves competitive multi-label classification while improving interpretability. In a reader study with expert radiologists, full reasoning traces increased confidence, supported error auditing, and reduced time to finalize reports. We release code and model NV-Reason-CXR-3B to facilitate community progress toward trustworthy, explainable AI in chest radiography and broader medical imaging tasks where reasoning quality is as critical as prediction quality.
Github Code: https://github.com/NVIDIA-Medtech/NV-Reason-CXR
Model: https://huggingface.co/nvidia/NV-Reason-CXR-3B
@@ 2.1 Human annotations
We used OpenAI’s whisper-large-v3 model to transcribe audio to text [7]. The system transcribes in the background; radiologists then review the transcript and edit if necessary. We observed frequent transcription errors, particularly for medical terminology. To mitigate this, we added a second pass using an LLM for domain-aware correction. We instructed the LLM that the text concerns chest X-ray radiology to bias corrections toward appropriate terms. This substantially reduced manual edits by radiologists, who previously had to fix spelling and terminology errors. For this correction stage we used the DeepSeek-V3-0324 LLM model [8].
@@ 2.5 Structured Report Query
INDICATION: Not provided. COMPARISON: None. TECHNIQUE: PA view of the chest. FINDINGS: Medical devices: Chest tube at right lung base. Central airways: Trachea and main bronchi patent. Lungs and pleural space: Small right pneumothorax; small right pleural effusion; parenchymal opacity in right lung; left lung clear. Mediastinum and hila: Mediastinum normal; hila normal; no adenopathy. Heart: Enlarged heart. Upper abdomen and diaphragm: No free subdiaphragmatic air; diaphragm normal. Bones and soft tissues: Normal. IMPRESSION: Small right-sided pneumothorax. Small right pleural effusion with parenchymal opacity. Enlarged heart
@@ 3.5 Verifiable Rewards for Chest X-Rays
The main reward function we use is abnormality set correctness, which is an intersection over union metric between predicted and ground-truth sets. Let $\hat{z}_{\text{cls}}=\Phi_{\text{cls}}(y)$ be the set extracted from $y$.
@@
Weights can upweight safety-critical findings (e.g., Lung Lesion), but in practice we used equal weights $w_{a}==1$. This reward function is by definition bound to (0, 1) interval.
@@ 3.6 Training Details
Backbone. We adopt Qwen2.5-VL-3B-Instruct [^1] https://hf.co/Qwen/Qwen2.5-VL-3B-Instruct, a vision–language model that combines a ViT-based vision encoder with a decoder-only LLM. Qwen2.5-VL features several architectural and training improvements, e.g. an efficient vision stack (windowed attention in ViT) and alignment to the language backbone with RMSNorm and SwiGLU—which together yield strong zero-/few-shot multimodal performance while remaining parameter-efficient for domain adaptation [15].
@@ 3.7 SFT
The training was performed on 4 nodes of 8xH100 NVIDIA GPU machines. All experiments use $bfloat16$ training with gradient checkpointing, gradient accumulation steps of 8. For GRPO each prompt has a number of generations equal to 16, with temperature of 1. Each gpu was set to handle 1 prompt. AdamW optimizer was used with cosine schedule starting from a learning rate of $1e^{-6}$. Grad norm clipping was set to 0.2. Optimization was run for 10 epochs.
@@ 3.9 Implementation details
Many design choices were inspired by Open-R1222https://github.com/huggingface/open-r1 - a great open source project by Hugginface aiming to reproduce DeepSeek-R1 [2] performance for text only inputs in PyTorch. Both SFT and GRPO training loops take advantage of several Hugginface libraries, including TRL, Transformers and Accelerate. We make our training code public [^3] https://github.com/NVIDIA-Medtech/NV-Reason-CXR.
@@ 4.1 Subjective Evaluation with Expert Radiologists
    After a short distraction, the case is re-presented with the abnormalities-only list (AI generated); the reader may revise findings/impressions and time.
@@
    After a second distraction, the case is shown with the full reasoning text and its structured report form. The reader can copy-paste the structured report, double check it and edit if necessary. The total time is recorded again.
@@ 4.2 Numeric Evaluation
To evaluate multi-label prediction, model outputs are taken from the structured <answer> block which are already mapped to 14 classes. MedGemini work considered both multi-select prompt (asking for all labels at once) and per-label Yes/No prompts, and found per-label prompts provided better macro-F1 metric scores. In our work we use a single multi-select prompt to be consistent with the overall model workflow. The prompt was "Find abnormalities and support devices", and the reasoning text in <think> tags was ignored.
@@
Training used data only from the MIMIC-CXR training set. The SFT stage used a subset of 100K cases (images + reports), whereas the RL stage used all 227K cases (images + report derived labels). OOD evaluation uses the official CheXpert test set.
@@
We report macro F1 scores. Clinical classification tasks are often class-imbalanced: positive (abnormal) cases are relatively rare, yet clinically critical. Metrics that are dominated by the majority class (e.g., accuracy) can therefore be misleading. The F1 score directly balances precision and recall (sensitivity), making it well-suited to summarize performance when both missed findings and false alarms have tangible clinical costs.